the length of a rectangle is 4 less than twice the width the perimeter of the rectangle is 198 find the dimensions of the rectangle

Answer by checkley77(12569)

L=2W-4
2L+2W=198
2(2W-4)+2W=198
4W-8+2W=198
6W=198+8
6W=206
W=206/6
W=34.3333 ANS. FOR THE WIDTH.
L=2*34.3333-4
L=68.6667-4
L=64.6667 ANS FOR THE LENGTH.
PROOF:
2*64.6667+2*34.3333=198
129.3333+68.6667=198
198=198